# Probing Top Quark - Electron Interactions at Future Colliders

### v1

@@ Abstract
Top quark interactions offer a window into possible new high scale physics and many models of new physics predict that the top quark interactions will deviate significantly from those predicted by the Standard Model (SM). We present an analysis of the experimental restrictions on anomalous 4-fermion $e^{+}e^{-}t{\overline{t}}$ operators that is accurate to next-to-leading order (NLO) in both the electroweak and QCD interactions within the Standard Model Effective Field Theory framework. At NLO, there is sensitivity to an extended set of anomalous interactions beyond those probed at leading order. A comparison of current limits from electroweak precision observables, along with expected future limits from Drell-Yan and $t{\overline{t}}e^{+}e^{-}$ production at the high luminosity LHC, from deep inelastic scattering at the EIC, and from projected sensitivities at the future FCC-ee and CEPC machines demonstrates that each of these programs extends the precision understanding of the interactions of top quarks.
@@ I Introduction and Motivation
The top quark plays a special role in the Standard Model. Due to its large mass, the couplings of the top quark to the Higgs boson and to longitudinal gauge bosons are much larger than those of the light quarks, suggesting that the top quark may be fundamentally different from the lighter fermions. Precision probes of top quark properties can thus help to elucidate questions of flavor in the quark sector and to determine whether there is new high scale physics affecting top quark interactions.
@@
Limits on heavy new physics are often expressed in the framework of effective field theories (EFTs). If the unknown new physics is much heavier than the weak scale, EFTs offer a consistent way to compare the capabilities of different current and future colliders. By combining measurements from many sources, more stringent bounds are often obtained than are possible from a single experiment. In this letter, we consider restrictions on operators involving the top quark in association with electrons from existing electroweak precision observables (EWPOs), neutral Drell-Yan and $pp\to t{\overline{t}}l^{+}l^{-}$ production at the high luminosity LHC (HL-LHC), polarized deep inelastic scattering (DIS) at the future Electron-Ion Collider (EIC)[1], and from future Z pole measurements, Higgs-Z associated production and top-pair production at the proposed FCC-ee and CEPC $e^{+}e^{-}$ colliders [2, 3, 4, 5]. For each of these experimental measurements, we present the limits on anomalous top quark-electron interactions and combine our results into a global fit [^1] We do not discuss the proposed ILC [50], as we have not optimized for polarization.. We assume that the only anomalous interactions are 4-fermion operators with top quarks and leptons, and set all other non-SM interactions to zero.
@@
Our study computes observables in the dimension-6 SMEFT and is accurate to next-to-leading order (NLO) in both the QCD and electroweak sectors. At NLO, there is sensitivity to new interactions that are not probed at lowest order (LO). Interactions with 4 light fermions are strongly restricted by current low energy data [7, 8]. The two-fermion top-quark operators related to the vertices $t\bar{t}V$, $tbW$, and $t\bar{t}H$, along with the four-fermion top-quark operators $q\bar{q}t\bar{t}$ are also stringently constrained [9, 10, 11, 12, 13, 14, 15]. In this letter, we focus on the two-lepton-two-top-quark operators, which are currently only weakly constrained by the current LHC $pp\to t\bar{t}\ell^{+}\ell^{-}$ measurements [16].
@@
One of the novel features of our study is that all relevant NLO electroweak corrections are included in the computations of observables. NLO QCD corrections can be automated [17, 18], but electroweak corrections must be computed on a case by case basis in the SMEFT. The NLO electroweak contributions to Drell-Yan and DIS processes in the dimension-6 SMEFT from 4-fermion top-electron operators are presented here for the first time.
@@ II SMEFT NLO Calculations
The calculation of physical observables is both an expansion in powers of $\frac{1}{\Lambda}$ and in loops, $\sim\frac{1}{16\pi^{2}}$ [19]. We present ${\cal{O}}(\frac{1}{16\pi^{2}\Lambda^{2}})$ NLO SMEFT results that are accurate to one loop in the QCD and electroweak interactions using the dimension-6 SMEFT. The NLO results include both virtual one-loop contributions and the real emission of photons and gluons and are performed with dimensional regularization in $D=4-2\epsilon$ dimensions. As input parameters, we take $M_{W}=80.369\pm.013~{}\rm GeV$, $M_{Z}=91.1876\pm.0021~{}\rm GeV$ , $G_{F}=1.1663787(6)\times 10^{-5}~{}/\rm GeV^{2}$, $M_{h}=125.10\pm 0.14~{}\rm GeV$, $M_{t}=172.9\pm 0.5~{}\rm GeV$, and $\alpha_{s}(M_{Z})=0.1181\pm 0.0011$. The CKM matrix is set to be diagonal and all fermions other than the top are assumed to be massless. We use a hybrid renormalization scheme: SM parameters are renormalized on-shell, while coefficient functions, $C_{i}^{d}$ are renormalized using ${\overline{\rm MS}}$ [20].
Some of the processes we consider have been previously computed at NLO in the SMEFT: the NLO EWPOs [21, 22, 23] and $e^{+}e^{-}\rightarrow Zh$ cross section [24, 25] are known analytically. The NLO results for all dimension-6 SMEFT contributions except from the 4-fermion interactions for Drell-Yan are known [26, 27], while the 4-fermion results for Drell-Yan and DIS are new in this work.
We use FeynRules [28] routines to convert SMEFT $R_{\xi}$ Feynman rules [29, 30] into a FeynArts [31] model file, then using FeynCalc [32], we compute helicity amplitudes and reduce $1$-loop integrals to Passarino-Veltman integrals which are evaluated using Package-X/FeynHelpers [33, 34, 35] and Looptools [36]. We do not fix the gauge in the SMEFT Lagrangian and we verify the cancellation of $\xi$ terms, in order to have a further check on our computations. To be consistent in the SMEFT expansion, our approach allows us to consider Feynman diagrams with at most a single SMEFT operator insertion. This ensures renormalizability by avoiding divergences that cannot be removed without introducing higher-dimension operators at leading order.
@@
$\mathcal{O}^{(3),1133}_{\ell q},\mathcal{O}^{(1),1133}_{\ell q},\mathcal{O}^{1133}_{\ell u},\mathcal{O}^{3311}_{qe},\mathcal{O}^{1133}_{eu},$ (2)
where the four indices in the superscript denote the flavor of the four fermions. The operators we consider are identical to those of the top-centric model of [22]. The operators in the Warsaw basis [37] are defined in Table 1 and they contribute to the effective interactions $e^{+}e^{-}q\bar{q}$ (where $q$ is a light quark) and $e^{+}e^{-}Zh$ at NLO. We do not consider the chirality flipped scalar and tensor operators as they do not interfere with the SM contribution in the limit of massless leptons. Some sample Feynman diagrams are shown in Fig. 1. The DIS process is obtained by rotating the Feynman diagram on the left-hand side by $90^{o}$’s.
@@ II.1 Drell-Yan
NLO QCD/electroweak results for neutral Drell-Yan production including all contributing operators except for four-fermion operators are given in [26]. The four-fermion operators involving light quarks contribute at tree level [7] and are stringently limited from existing Drell-Yan measurements [38, 39, 40]. The four-fermion operators involving electrons and top quarks defined in eq. (2) contribute at NLO and analytic results for the helicity amplitudes are presented here for the first time and documented in the supplemental material. In the limit where the partonic center of mass energy ${\hat{s}}\gg m_{t}^{2}$, the NLO SMEFT contributions to the spin- and color-averaged partonic cross sections, ${\hat{\sigma}}^{\rm NLO}_{\rm DY}$, can be simplified,
@@
$\bar{u}u\to e^{+}e^{-}$ $(X_{0,i})$ $\frac{w(2w+z)}{4}$ $-\frac{14w^{2}-31wz+17z^{2}}{36}$ $-\frac{14w^{2}-31wz+17z^{2}}{18}$ $-\frac{17(w-z)^{2}}{18}$ $-\frac{17(w-z)^{2}}{9}$
$(X_{1,i})$ $\frac{3(88w^{2}-164wz+85z^{2})}{16}$ $-\frac{14w^{2}-31wz+17z^{2}}{2}$ $-\frac{32w^{2}-22wz+17z^{2}}{8}$ $-17(w-z)^{2}$ $-\frac{8w^{2}-25wz+17z^{2}}{4}$
$\bar{d}d\to e^{+}e^{-}$ $(X_{0,i})$ $\frac{w(4w-z)}{4}$ $\frac{13wz-8w^{2}-5z^{2}}{36}$ $\frac{13wz-8w^{2}-5z^{2}}{18}$ $-\frac{5(w-z)^{2}}{18}$ $-\frac{5(w-z)^{2}}{9}$
$(X_{1,i})$ $\frac{3(40w^{2}-56wz+25z^{2})}{16}$ $-\frac{8w^{2}-13wz+5z^{2}}{2}$ $-\frac{44w^{2}-22wz+5z^{2}}{8}$ $-5(w-z)^{2}$ $-\frac{14w^{2}-19wz+5z^{2}}{4}$
@@
The logarithmic terms are sub-dominant at the LHC due to the suppression of $(m_{t}^{2}/{\hat{s}})$ or $(m_{t}^{2}/{\hat{s}})^{2}$, and this is clear in Fig. 2, where we show the partonic SM cross sections (black curve) and the contributions from a representative SMEFT operator (solid red curve) as a function of $\sqrt{\hat{s}}=m_{\ell\ell}$. The contributions from the logarithmic and non-logarithmic terms are shown by the dashed and dotted red curves, respectively. For illustration, we set $C^{(3),1133}_{\ell q}=1$, and the remaining coefficients to zero. Similar features hold for the other operators. The overall positive (negative) sign in $X_{0,i}$ indicates a constructive (destructive) interference with the SM.
@@ II.2 Electron Ion Collider (EIC)
By interchanging the Mandelstam variables, ${\hat{s}}\leftrightarrow{\hat{t}}$, the NLO results for polarized electron-quark scattering at the future EIC are obtained from the Drell-Yan results. The EIC is able to produce an electron beam with $P_{e}=70\%$ polarization and a systematic uncertainty $\lesssim$1[1] at $\sqrt{s}=140$ GeV. To cancel the large SM photon contributions, we compute the left-right asymmetry
@@
as our observable at the EIC, where $\tilde{\sigma}_{L(R)}\equiv P_{e}\sigma_{L(R)}+(1-P_{e})\sigma_{R(L)}$, and $\sigma_{L(R)}$ is the hadronic cross section for the deep-inelastic scattering of a left (right)-handed electron.
@@ II.3 Electroweak Precision Observables
The NLO SMEFT results for Z-pole observables were first computed in [21], and the results with an arbitrary flavor structure presented in [22, 23]. We include the following Z-pole observables,
@@
II.4 Higgstrahlung and top pair production at FCC-ee
The four-fermion electron-top operators contribute to Higgstrahlung at NLO as depicted by the middle panel in Fig. 1, and the complete NLO SMEFT result for $\sigma(e^{+}e^{-}\rightarrow Zh)$ is  [24, 25],
$\sigma_{e^{+}e^{-}\rightarrow Zh}=\sigma_{e^{+}e^{-}\rightarrow Zh}^{\rm NLO,SM}\biggl{[}1+\sum_{i}\biggl{(}\Delta_{i}+{\overline{\Delta_{i}}}\log\biggl{(}\frac{\mu^{2}}{\Lambda^{2}}\biggr{)}\biggr{)}\biggr{]}\,,$ (6)
Numerical values for $\Delta_{i}$ and ${\overline{\Delta}}_{i}$ are in Table 3. The relevant SM cross sections are $\sigma_{e^{+}e^{-}\rightarrow Zh}^{\rm NLO,SM}(\sqrt{s}=240~{}{\rm GeV})=232~{}{\rm fb}$ and $\sigma_{e^{+}e^{-}\rightarrow Zh}^{\rm NLO,SM}(\sqrt{s}=365~{}{\rm GeV})=113~{}{\rm fb}$ [41] . The projected precision on the Higgstrahlung cross section at FCC-ee is $0.52\%$ at $\sqrt{s}=240~{}{\rm GeV}$ and $1.35\%$ at $\sqrt{s}=365~{}{\rm GeV}$[42].
Table 3: NLO electroweak contributions to the total cross section for $e^{+}e^{-}\rightarrow Zh$ from anomalous top quark- electron interactions as defined in eq. (6). We set $\Lambda=1~{}{\textrm{TeV}}$.
$\sqrt{s}=240~{}{\rm GeV}$ $\sqrt{s}=365~{}{\rm GeV}$
   $\Delta_{i}/\Lambda^{2}$    $\bar{\Delta}_{i}/\Lambda^{2}$   $\Delta_{i}/\Lambda^{2}$   $\bar{\Delta}_{i}/\Lambda^{2}$
$C_{eu}^{1133}$ $\hphantom{\textrm{-}}0.01\cdot 10^{\textrm{-}3}$$\textrm{-}1.39\cdot 10^{\textrm{-}2}$$\textrm{-}3.73\cdot 10^{\textrm{-}2}$$\textrm{-}3.23\cdot 10^{\textrm{-}2}$
$C_{lu}^{1133}$$\textrm{-}0.02\cdot 10^{\textrm{-}3}$ $\hphantom{\textrm{-}}1.73\cdot 10^{\textrm{-}2}$ $\hphantom{\textrm{-}}4.64\cdot 10^{\textrm{-}2}$ $\hphantom{\textrm{-}}4.01\cdot 10^{\textrm{-}2}$
$C_{lq}^{(1)1133}$$\textrm{-}0.37\cdot 10^{\textrm{-}2}$$\textrm{-}1.80\cdot 10^{\textrm{-}2}$$\textrm{-}6.09\cdot 10^{\textrm{-}2}$$\textrm{-}4.18\cdot 10^{\textrm{-}2}$
$C_{lq}^{(3)1,1,3,3}$$\textrm{-}0.37\cdot 10^{\textrm{-}2}$ $\hphantom{\textrm{-}}1.29\cdot 10^{\textrm{-}2}$ $\hphantom{\textrm{-}}4.54\cdot 10^{\textrm{-}2}$ $\hphantom{\textrm{-}}3.29\cdot 10^{\textrm{-}2}$
$C_{qe}^{3311}$ $\hphantom{\textrm{-}}0.30\cdot 10^{\textrm{-}2}$ $\hphantom{\textrm{-}}1.45\cdot 10^{\textrm{-}2}$ $\hphantom{\textrm{-}}4.90\cdot 10^{\textrm{-}2}$ $\hphantom{\textrm{-}}3.36\cdot 10^{\textrm{-}2}$
At $\sqrt{s}=365$ GeV, the process $e^{+}e^{-}\rightarrow t{\overline{t}}$ becomes accessible and the electron-top four-fermion operators can be probed directly at LO as shown in the right panel of Fig. 1. The tree level result is

### v2

@@ Abstract
Top quark interactions offer a window into possible new high scale physics and many models of new physics predict that the top quark interactions will deviate significantly from those predicted by the Standard Model. We present an analysis of the experimental restrictions on anomalous 4-fermion $e^{+}e^{-}t{\overline{t}}$ operators that is accurate to next-to-leading order (NLO) in both the electroweak and QCD interactions within the Standard Model Effective Field Theory framework. At NLO, there is sensitivity to an extended set of anomalous interactions beyond those probed at leading order. A comparison of current limits from electroweak precision observables, along with expected future limits from Drell-Yan and $t{\overline{t}}e^{+}e^{-}$ production at the high luminosity LHC, from deep inelastic scattering at the EIC, and from projected sensitivities at the future FCC-ee and CEPC machines demonstrates that each of these programs extends the precision understanding of the interactions of top quarks.
preprint: IFT-UAM/CSIC-25-74
@@ I Introduction and Motivation
The top quark plays a special role in the Standard Model (SM). Because of its large mass, the couplings of the top quark to the Higgs boson and to longitudinal gauge bosons are much larger than those of the light quarks, suggesting that the top quark may be fundamentally different from the lighter fermions. Precision probes of top quark properties can thus help to elucidate questions of flavor in the quark sector and to determine whether there is new high scale physics affecting top quark interactions.
@@
Limits on heavy new physics are often expressed in the framework of effective field theories (EFTs). If the unknown new physics is much heavier than the weak scale, EFTs offer a consistent way to compare the capabilities of different current and future colliders. By combining measurements from many sources, more stringent bounds are often obtained than are possible from a single experiment. In this letter, we consider restrictions on operators involving the top quark in association with electrons from existing electroweak precision observables (EWPOs), neutral Drell-Yan and $pp\to t{\overline{t}}l^{+}l^{-}$ production at the high luminosity LHC (HL-LHC), polarized deep inelastic scattering (DIS) at the future Electron-Ion Collider (EIC)[AbdulKhalek:2021gbh], and from future $Z$-pole measurements, Higgs-Z associated production and top-pair production at the proposed FCC-ee and CEPC $e^{+}e^{-}$ colliders [FCC:2025lpp, Antusch:2025lpm, Altmann:2025feg, Celada:2024mcf]. For each of these experimental measurements, we present the limits on anomalous top quark-electron interactions and combine our results into a global fit [^1] We do not discuss the proposed ILC [LinearColliderVision:2025hlt], as we have not optimized for polarization.. We assume that the only anomalous interactions are 4-fermion operators with top quarks and leptons, and set all other non-SM interactions to zero.
@@
Our study computes observables in the dimension-6 Standard Model Effective Field Theory (SMEFT) and is accurate to next-to-leading order (NLO) in both the QCD and electroweak sectors. At NLO, there is sensitivity to new interactions that are not probed at lowest order (LO). Interactions with 4 light fermions are strongly restricted by current low energy data [deBlas:2013qqa, Falkowski:2017pss]. The two-fermion top-quark operators related to the vertices $t\bar{t}V$, $tbW$, and $t\bar{t}H$, along with the four-fermion top-quark operators $q\bar{q}t\bar{t}$ are also stringently constrained [Durieux:2018tev, Durieux:2018ggn, Durieux:2019rbz, Liu:2022vgo, Durieux:2022cvf, Severi:2022qjy, Cornet-Gomez:2025jot]. In this letter, we focus on the two-lepton-two-top-quark operators, which are currently only weakly constrained by the current LHC $pp\to t\bar{t}\ell^{+}\ell^{-}$ measurements [CMS:2023xyc, ATLAS:2025yww].
@@
One of the novel features of our study is that all relevant NLO electroweak corrections are included in the computations of observables. NLO QCD corrections can be automated [Brivio:2020onw, Degrande:2020evl], but electroweak corrections must be computed on a case by case basis in the SMEFT. The NLO electroweak contributions to Drell-Yan and DIS processes in the dimension-6 SMEFT from 4-fermion top-electron operators are presented here for the first time.
@@ II SMEFT NLO Calculations
The calculation of physical observables is both an expansion in powers of $\frac{1}{\Lambda}$ and in loops, $\sim\frac{1}{16\pi^{2}}$ [Buchalla:2022vjp]. We present ${\cal{O}}(\frac{1}{16\pi^{2}\Lambda^{2}})$ NLO SMEFT results that are accurate to one loop in the QCD and electroweak interactions using the dimension-6 SMEFT. The NLO results include both virtual one-loop contributions and the real emission of photons and gluons and are performed with dimensional regularization in $D=4-2\epsilon$ dimensions. As input parameters, we take $M_{W}=80.369\pm.013~\rm GeV$, $M_{Z}=91.1876\pm.0021~\rm GeV$ , $G_{F}=1.1663787(6)\times 10^{-5}~/\rm GeV^{2}$, $M_{h}=125.10\pm 0.14~\rm GeV$, $M_{t}=172.9\pm 0.5~\rm GeV$, and $\alpha_{s}(M_{Z})=0.1181\pm 0.0011$. The CKM matrix is set to be diagonal and all fermions other than the top are assumed to be massless. We use a hybrid renormalization scheme: SM parameters are renormalized on-shell, while coefficient functions, $C_{i}^{d}$ are renormalized using $\overline{\rm MS}$} [Alonso:2013hga].
Some of the processes we consider have been previously computed at NLO in the SMEFT: the NLO EWPOs [Dawson:2019clf, Bellafronte:2023amz, Biekotter:2025nln] and $e^{+}e^{-}\rightarrow Zh$ cross section [Asteriadis:2024xuk, Asteriadis:2024xts] are known analytically. The NLO results for all dimension-6 SMEFT contributions except from the 4-fermion interactions for Drell-Yan are known [Dawson:2021ofa, Dawson:2018dxp], while the 4-fermion results for Drell-Yan and DIS are new in this work.
We use FeynRules [feynrul] routines to convert SMEFT $R_{\xi}$ Feynman rules [Dedes:2017zog, Dedes:2023zws] into a FeynArts [feynarts] model file, then using FeynCalc [feync], we compute helicity amplitudes and reduce $1$-loop integrals to Passarino-Veltman integrals which are evaluated using Package-X/FeynHelpers [Patel:2015tea, Patel:2016fam, Shtabovenko:2016whf] and Looptools [Hahn:1998yk]. We do not fix the gauge in the SMEFT Lagrangian and we verify the cancellation of $\xi$ terms, in order to have a further check on our computations. To be consistent in the SMEFT expansion, our approach allows us to consider Feynman diagrams with at most a single SMEFT operator insertion. This ensures renormalizability by avoiding divergences that cannot be removed without introducing higher-dimension operators at leading order.
@@
$\mathcal{O}^{(3),1133}_{\ell q},\mathcal{O}^{(1),1133}_{\ell q},\mathcal{O}^{3311}_{qe},\mathcal{O}^{1133}_{\ell u},\mathcal{O}^{1133}_{eu},$ (2)
where the four indices in the superscript denote the flavor of the four fermions [Aguilar-Saavedra:2010uur]. Please note that we still refer to it as an NLO electroweak correction when a top quark appears in the loop, even though such contributions are leading order for the corresponding operators. The operators we consider are identical to those of the top-centric model of [Bellafronte:2023amz]. The operators in the Warsaw basis [Grzadkowski:2010es] are defined in Table 1 and they contribute to the effective interactions $e^{+}e^{-}q\bar{q}$ (where $q$ is a light quark) and $e^{+}e^{-}Zh$ at NLO. We do not consider the chirality flipped scalar and tensor operators as they do not interfere with the SM contribution in the limit of massless leptons. Nevertheless, those operators can be important in some precise measurements of the SM suppressed processes [Vasquez:2019muw]. Some sample Feynman diagrams are shown in Fig. 1. The DIS process is obtained by rotating the Feynman diagram on the left-hand side by $90^{o}$.
@@ II.1 Drell-Yan
NLO QCD/electroweak results for neutral Drell-Yan production including all contributing operators except for four-fermion operators are given in [Dawson:2021ofa]. The four-fermion operators involving light quarks contribute at tree level [deBlas:2013qqa] and are stringently limited from existing Drell-Yan measurements [Farina:2016rws, Panico:2021vav, Torre:2020aiz]. The four-fermion operators involving electrons and top quarks defined in eq. (2) contribute at NLO and analytic results for the helicity amplitudes are presented here for the first time and documented in the Supplemental Material [appendix]. In the limit where the partonic center of mass energy ${\hat{s}}\gg m_{t}^{2}$, the NLO SMEFT contributions to the spin- and color-averaged partonic cross sections, ${\hat{\sigma}}^{\rm NLO}_{\rm DY}$, can be simplified,
@@
$\bar{u}u\to e^{+}e^{-}$ $(X_{0,i})$ $\frac{w(2w+z)}{4}$ $\frac{14w^{2}-31wz+17z^{2}}{36}$ $\frac{14w^{2}-31wz+17z^{2}}{18}$ $\frac{17(w-z)^{2}}{18}$ $\frac{17(w-z)^{2}}{9}$
$(X_{1,i})$ $\frac{3(88w^{2}-164wz+85z^{2})}{16}$ $\frac{14w^{2}-31wz+17z^{2}}{2}$ $\frac{32w^{2}-22wz+17z^{2}}{8}$ $17(w-z)^{2}$ $\frac{8w^{2}-25wz+17z^{2}}{4}$
$\bar{d}d\to e^{+}e^{-}$ $(X_{0,i})$ $\frac{w(4w-z)}{4}$ $\frac{13wz-8w^{2}-5z^{2}}{36}$ $\frac{13wz-8w^{2}-5z^{2}}{18}$ $\frac{5(w-z)^{2}}{18}$ $\frac{5(w-z)^{2}}{9}$
$(X_{1,i})$ $\frac{3(40w^{2}-56wz+25z^{2})}{16}$ $\frac{8w^{2}-13wz+5z^{2}}{2}$ $\frac{44w^{2}-22wz+5z^{2}}{8}$ $5(w-z)^{2}$ $\frac{14w^{2}-19wz+5z^{2}}{4}$
@@
[…]
@@ II.2 Electron Ion Collider (EIC)
By interchanging the Mandelstam variables, $^{\hat{s}}\leftrightarrow{\hat{t}}$, the NLO results for polarized electron-quark scattering at the future EIC are obtained from the Drell-Yan results. The EIC is able to produce an electron beam with $P_{e}=70\%$ polarization and a systematic uncertainty $\lesssim$1% [AbdulKhalek:2021gbh] at $\sqrt{s}=140$ GeV. To cancel the large SM photon contributions, we compute the left-right asymmetry
@@
as our observable at the EIC, where $\tilde{\sigma}_{L(R)}\equiv\sigma_{L(R)}(1+P_{e})/2+\sigma_{R(L)}(1-P_{e})/2$, and $\sigma_{L(R)}$ is the hadronic cross section for the deep-inelastic scattering of a left (right)-handed electron.
@@ II.3 Electroweak Precision Observables
The NLO SMEFT results for $Z$-pole observables were first computed in [Dawson:2019clf], and the results with an arbitrary flavor structure presented in [Bellafronte:2023amz, Biekotter:2025nln]. We include the following $Z$-pole observables,
@@
II.4 FCC-ee
The next run above the $Z$-pole at the FCC-ee is at the $WW$ threshold ($\sqrt{s}=162$ GeV). The first three operators listed in eq. (2) involving the left-handed quark doublet can modify the cross section of $b{\overline{b}}$ pair production. At LO at $\sqrt{s}=162$ GeV,
$\sigma_{e^{-}e^{+}\to b\bar{b}}~(\rm pb)=5.97-\biggl\{0.69C^{3311}_{qe}$
$-1.96(C^{(3),1133}_{\ell q}+C^{(1),1133}_{\ell q})\biggr\}\biggl(\frac{1~\textrm{TeV}}{\Lambda}\biggr)^{2}.$ (6)
The FCC-ee at $\sqrt{s}=162$ GeV is expected to reach a precision of $2\times 10^{-4}$ on $\Delta R_{b}/R_{b}$ [Greljo:2024ytg].
The four-fermion electron-top operators contribute to Higgstrahlung at NLO as depicted by the middle panel in Fig. 1, and the complete NLO SMEFT result for $\sigma(e^{+}e^{-}\rightarrow Zh)$ can be found in  [Asteriadis:2024xuk, Asteriadis:2024xts].
At $\sqrt{s}=365$ GeV, the process $e^{+}e^{-}\rightarrow t{\overline{t}}$ becomes accessible and the electron-top four-fermion operators can be probed directly at LO as shown in the right panel of Fig. 1. The tree-level result at $\sqrt{s}=365$ GeV is